# Dance to Demise – How Massive Stars May Form Dense Circumstellar Shells Before Explosion

### v1

@@ Abstract
We investigate the evolution of red supergiant (RSG) progenitors of core-collapse (CC) supernovae (SNe) with initial masses between 122012-20 M focusing on effects of enhanced mass loss due to pulsation-driven instabilities in their envelopes and subsequent dynamical ejections during advanced stages of nuclear burning. Using time-dependent mass loss rates from detailed MESA stellar evolution models, including prescriptions for both pulsation-driven superwinds and shock-induced ejections, we construct the circumstellar medium (CSM) before the SN explosion. We calculate resulting CSM density profiles and column densities considering the radiation-driven acceleration of the stellar wind. Our models produce episodes of enhanced mass loss ($\sim 10^{-4}-10^{-2}$ M $\rm{yr}^{-1}$) in the last centuries-decades before explosion forming dense CSM ($\gtrsim 10^{-15}~\rm{g~cm}^{-3}$ at distances $\lesssim 10^{15}~\rm{cm}$) – consistent with multi-wavelength observations of Type II SNe such as SN 2023ixf, SN 2020ywx, SN 2017hcc, SN 2005ip and SN 1998S. The formation of such dense CS shells, as predicted by our single star RSG models, provides a natural explanation for observed flash-ionization signatures, X-ray and radio emission, and has important implications for dust formation around Type II SNe.
@@ 1 Background
However, several factors complicate the task of deriving reliable mass loss rates ($\dot{M}$) for cool, possibly dust-enshrouded, red supergiant (RSG) progenitors (J. T. van Loon et al., 2005) of Type II SNe (L. Decin, 2021). At the same time, there is a growing consensus that a high fraction of Type II SNe are preceded by phases of heavy mass loss (K.-R. Hinds et al., 2025) whose signatures can be seen from the post-explosion shock interaction with the circumstellar medium (CSM) created by the wind of the SN progenitor (P. Chandra, 2018; L. Dessart, 2024; V. V. Dwarkadas, 2025). The circumstellar densities inferred from flash-ionization spectroscopy, X-ray and radio observations of SNe like SN 1998S (N. Chugai et al., 2002), SN 2010jl (C. Fransson et al., 2014), and SN 2023ixf (K. A. Chugai et al., 2023) point toward mass loss rates of $\sim 10^{-3}-10^{-1}$ M $\rm{yr}^{-1}$ within $\sim 10^{15}-10^{16}$ cm of the progenitor —- values that far exceed those from standard steady winds. The presence of narrow P Cygni absorption features, light curve plateaus with CSM interaction signatures, and late-time IR emission all corroborate the scenario of enhanced pre-SN mass loss. All these observations of post-explosion features are clearly at odds with latest mass loss prescriptions inferred from large samples of observed RSGs (E. R. Beasor et al., 2023; P. Massey et al., 2023; M. Yang et al., 2023; K. Antoniadis et al., 2024; L. Decin et al., 2024) as well as empirical mass loss rates derived from much smaller samples of stars (C. de Jager et al., 1988; H. Nieuwenhuijzen & C. de Jager, 1990) that are still used in state-of-the-art stellar evolution codes to model the RSG phase of massive stars (N. Mauron & E. Josselin, 2011; P. Massey et al., 2021).
@@ 2 Rationale of this study
The motivation behind this work is to address (a) whether a single star may incur heavy mass loss in the final stages before it explodes, that can be correlated with the presence of a dense CSM near the star, and (b) how the progenitor mass impacts the mass loss rates, CSM densities, and wind velocities. In this regard, our study accounts for progenitors of Type II supernovae, and we compare the CSM profiles of well-observed interacting SNe such as SN 2023ixf (J. E. Jencson et al., 2023; W. V. Jacobson-Galán et al., 2023; E. A. Zimmerman et al., 2024; N. Smith et al., 2023; K. A. Bostroem et al., 2024; N. A. J. et al., 2025), SN 2005ip (N. Smith et al., 2009; S. Katsuda et al., 2014), SN 1998S (N. Chugai et al., 2002; M. Pozzo et al., 2004; C. Fransson et al., 2005; J. Mauerhan & N. Smith, 2012; I. Shivvers et al., 2015), SN 2017hcc (P. Chandra et al., 2022), SN 2020ywx (R. Baer-Way et al., 2025). We argue that for single stars, enhanced pulsation near the end of their lives provides conditions favorable for escalated mass loss (S.-C. Yoon & M. Cantiello, 2010; Moriya, Takashi J. & Langer, Norbert, 2015). To demonstrate this effect, we simulate the evolution of massive stars with initial masses between $12$ and 2020 M, from their main-sequence until the end of neon burning (or until end of core O-burning, in some cases) using the MESA code, applying advanced mass loss prescriptions to incorporate effects of radial pulsations in the outer envelope of RSGs (Section 4).
@@ 3 Mass loss in RSGs
Mass loss from RSGs is a fundamental yet still uncertain ingredient in the late stages of stellar evolution (R. M. Humphreys, 2025; S. D. Van Dyk, 2025; J. T. van Loon, 2025; S. Ekström & C. Georgy, 2025). The physical mechanisms responsible for this process directly shape the CSM, influencing both the pre-SN environment and the post-explosion observables, including light curves, spectra, and dust formation. Observations of Type II SNe with dense CSM (e.g., SN 2009ip, SN 2010jl, SN 2023ixf and so on) have revealed that enhanced and often eruptive mass loss can occur shortly before CC (T. Zhang et al., 2012; E. O. Ofek et al., 2013; K. A. Chugai et al., 2023). In this section, we describe potential driving mechanisms for steady line-driven winds, pulsation-driven outflows including dynamical ejections and/or eruptions, turbulent pressure and convective shocks creating extended atmosphere/chromosphere for dust-driven winds, and enhanced mass loss due to binary interaction for RSGs.
@@ 3.1 Line-driven wind
Line-driven winds are driven by the transfer of momentum from photons to metal lines in the stellar atmosphere. This mechanism is efficient in hot, compact stars e.g., OB stars (J. S. Vink, 2017), but is considerably less effective in RSGs due to their cooler photospheres ($T_{\mathrm{eff}}\lesssim 4000$ K), where fewer atoms are ionized and UV photons are scarce (J. I. Castor et al., 1975; L. B. van den Hoek & M. A. T. Groenewegen, 1997). Typical mass loss rates from line-driven winds in RSGs range from $\sim 10^{-7}$ to $10^{-6}$ M $\rm{yr}^{-1}$ (E. R. Beasor et al., 2021; J. Merritt et al., 2025), which is inadequate to explain the dense CSM observed in the majority (upto $80\%$) of all Type II SNe wih early-time observations (V. Morozova et al., 2018; F. Förster et al., 2018).
@@ 3.2 Pulsation driven mass loss
Many RSGs exhibit large-amplitude radial pulsations with periods of several hundred days (M. D. Soraisam et al., 2018; Y. Ren et al., 2019), as observed in long-period variables (LPVs). Such radial motions in the H-rich envelope of RSGs are driven by the so-called $\kappa$-mechanism (N. Baker & R. Kippenhahn, 1965; T. Gastine & B. Dintrans, 2008) - an interplay between excess radiation pressure and gravitational pull on the outer layers of the RSG envelope. These pulsations can grow in amplitude to produce shocks at the stellar surface that levitate material above the photosphere, reducing the effective gravity and allowing other processes (e.g., radiation pressure on dust and molecules) to accelerate the material outward. However, the exact mechanism of mass loss due to the complex interplay of pulsations and convection in RSG envelopes remains ill-understood (T. Gastine & B. Dintrans, 2010), and attempts to reduce these inherently three-dimensional phenomena into one-dimensional prescriptions for use in evolutionary codes remain limited (S.-C. Yoon & M. Cantiello, 2010; Moriya, Takashi J. & Langer, Norbert, 2015).
@@ 3.2.1 Dynamical Ejections
In the advanced nuclear burning stages, vigorous convection and shell flashes can lead to hydrodynamic instabilities and formations of shock waves in the outer envelopes of RSGs. If these internal shocks propagate outward and reach the stellar surface, they can trigger impulsive ejections of envelope material. These “dynamical ejections” are thought to be responsible for the formation of dense, compact shells detected in SN progenitors such as SN 2009ip (J. C. Mauerhan et al., 2013; N. Smith, 2014). Simulations suggest that mass loss rates in such events can temporarily exceed $10^{-2}$ M $\rm{yr}^{-1}$, lasting sometimes for days to weeks or even sustain repeatedly for centuries (M. Clayton, 2018).
@@ 3.2.2 Eruptive mass loss
Eruptive mass loss may occur through surface instabilities driven by opacity changes, pulsationally-enhanced convective transport, wave-driven energy deposition or super-Eddington luminosities in RSG envelopes (E. Quataert & J. Shiode, 2012; J. Fuller, 2017; S. J. Cheng et al., 2024). These eruptions resemble luminous blue variable (LBV) behavior, though they are seen in cooler RSGs as well. The pre-SN outbursts seen in light echoes (e.g., SN 2013cu) are likely signatures of such events, with inferred $\dot{M}\sim 10^{-4}$$10^{-2}$ M $\rm{yr}^{-1}$ over timescales of months to years (G. Gräfener & J. S. Vink, 2016). Furthermore, statistical analysis of luminous stars in the Milky Way (C. S. Kochanek, 2011) and the study of $\eta$ Carinae analogs in nearby galaxies (R. Khan et al., 2015) reveal the presence of massive shells which hint towards instabilities triggered at the onset of C-burning leading to eruptions (R. Margutti et al., 2017).
@@ 4.1 Pulsation-Driven Superwind (PDSW)
Red Supergiants (RSGs) often exhibit long-period variability (M. D. Soraisam et al., 2018; Y. Ren et al., 2019; B. Hsu et al., 2024) that has been associated with changes in stellar radii due to excitation of different modes of pulsational instabilities through the $\kappa$-mechanism (H. Saio et al., 2016). Theoretically, A. Heger et al. (1997) demonstrated the growth of such pulsational instabilities in hydrodynamical evolutionary calculations and found that the growth rates in their non-linear numerical calculations to be proportional to the luminosity ($\mathrm{L}$) to mass ($\mathrm{M}$) ratio. S.-C. Yoon & M. Cantiello (2010) showed that the growth rate ($\eta$) of the amplitude of the surface velocities increases with $L/M$ ratio and/or decreasing thermal (Kelvin-Helmholtz) time scale of the envelope ($\tau_{KH,env}$), according to the relation:
@@
where $R$ is the stellar radius, $\tau_{KH,env}:=GMM_{env}/LR$ ($M_{env}$ is the mass of the H-rich envelope ) and $C_{1}=9.219\times 10^{-6},~C_{2}=0.0393$ which are calibrated against their hydrodynamic models. Figure 1 shows a linear least-squares fit for the relation given by Equation 1 for our hydrodynamic MESA model sequences with initial masses between 121812-18 M, which gives $C_{1}=9.393\times 10^{-6},C_{2}=0.0797$, in close agreement with the values of S.-C. Yoon & M. Cantiello (2010).
Figure 1: Growth rate of radial oscillations, $\eta$, as a function of stellar parameters (radius, mass, and Kelvin-Helmholtz time-scale of the envelope); the solid line represents a linear least-squares fit
Figure 2: (Top) Evolution of the stellar radius (black curve plotted on the left axis) in a set of hydrodynamic MESA model sequences with initial mass $M_{\mathrm{init}}=15~\mathrm{M}_{\odot}$ in the RSG phase until shock breakout at the surface when Mach Number ($v_{surf}/c_{sound}$) exceeds unity (shown in red on the right axis). (Bottom) Corresponding evolution of a set of MESA models with varying initial masses between $12-18\ M_{\odot}$ in the HR diagram
The top panel of Figure 2 shows a typical hydrodynamical model for a MESA run with initial mass $M_{\mathrm{init}}=15$ M illustrating the growth of the radial pulsations until shock break-out at the surface where the Mach number reaches unity. The bottom panel of Figure 2 shows the corresponding evolution of a set of MESA runs with varying initial masses between 121812-18 M in the Hertzspring-Russell (HR) diagram. Capturing these radial pulsations with the typical time period, $P\sim 2-5~yrs$, requires using much smaller time steps, $\delta t\leq 0.01~\text{years}$ (B. Paxton et al., 2013), that makes modeling such pulsations a challenging task for evolutionary codes which typically use much larger time steps during the onset and growth of these pulsations. We do not attempt to resolve these pulsations for our full evolutionary runs (from pre main-sequence until core oxygen-depletion) for computational feasibility but incorporate potential impact of the growth of such pulsations observed in advanced phases (beyond core-helium burning) on the mass loss history of our stellar models. To do this, we evolve a set of hydrostatic MESA models until end of core-helium burning using the standard Dutch wind scheme in MESA. Thereafter, each run is continued with the hydrodynamic term on and replacing the wind mass loss rates with the ”superwind” (SW) mass loss prescription of S.-C. Yoon & M. Cantiello (2010) given by
@@
for $\eta>1$ using the growth rate expression as given in Equation 1 with our calibrated constants. Here, $\dot{M}_{\mathrm{JNH88}}$ is the mass loss rate from de C. de Jager et al. (1988) scaled down by a factor of $0.2$ for reasons explained in Section 4, and $\alpha$ is a parameter which sets the strength and consequently the duration of the SW-phase (Section 4.2). Unlike S.-C. Yoon & M. Cantiello (2010) who used a fixed $\alpha~=~1.75$, we conduct a thorough exploration of the dependence of the SW on both the initial mass $\mathrm{M}_{init}=12-20$ M and value of $\alpha=0-5$.

### v2

@@ Abstract
We investigate the evolution of red supergiant (RSG) progenitors of core-collapse (CC) supernovae (SNe) with initial masses between 122012-20 M focusing on the effects of enhanced mass loss due to pulsation-driven instabilities in their envelopes and subsequent dynamical ejections during advanced stages of nuclear burning. Using time-dependent mass loss from detailed MESA stellar evolution models, including a parameterized prescription for pulsation-driven superwinds and time-averaged mass loss rates attributed to resulting shock-induced ejections, we construct the circumstellar medium (CSM) before the SN explosion. We calculate resulting CSM density profiles and column densities considering the acceleration of the stellar wind. Our models produce episodes of enhanced mass loss ($\sim 10^{-4}-10^{-2}$ M $\rm{yr}^{-1}$) in the last centuries-decades before explosion forming dense CSM ($\gtrsim 10^{-15}~\rm{g~cm}^{-3}$ at distances $\lesssim 10^{15}~\rm{cm}$) – consistent with those inferred from multi-wavelength observations of Type II SNe such as SN 2023ixf, SN 2020ywx, SN 2017hcc, SN 2005ip and SN 1998S. The formation of such dense CS shells, within the explored range of our single star RSG models, provides a natural explanation for observed flash-ionization signatures, X-ray and radio emission, and has important implications for dust formation around Type II SNe.
@@ 1 Background
However, several factors complicate the task of deriving reliable mass loss rates ($\dot{M}$) for cool, possibly dust-enshrouded, red supergiant (RSG) progenitors of Type II SNe (J. T. van Loon et al., 2005; L. Decin, 2021). At the same time, there is a growing consensus that a high fraction of Type II SNe are preceded by phases of heavy mass loss (K.-R. Hinds et al., 2025) whose signatures can be seen from the post-explosion shock interaction with the circumstellar medium (CSM) created by the wind of the SN progenitor (P. Chandra, 2018; L. Dessart, 2024; V. V. Dwarkadas, 2025). The circumstellar densities inferred from flash-ionization spectroscopy, X-ray and radio observations of SNe like SN 1998S (N. Chugai et al., 2002), SN 2010jl (C. Fransson et al., 2014), and SN 2023ixf (K. A. Chugai et al., 2023) point toward mass loss rates of $\sim 10^{-3}-10^{-1}$ M $\rm{yr}^{-1}$ within $\sim 10^{15}-10^{16}$ cm of the progenitor —- values that far exceed those from standard steady winds. The presence of narrow P Cygni absorption features, light curve plateaus with CSM interaction signatures, and late-time IR emission all corroborate the scenario of enhanced pre-SN mass loss. All these observations of post-explosion features are clearly at odds with latest mass loss prescriptions inferred from large samples of observed RSGs (E. R. Beasor et al., 2023; P. Massey et al., 2023; M. Yang et al., 2023; K. Antoniadis et al., 2024; L. Decin et al., 2024) as well as empirical mass loss rates derived from much smaller samples of stars (C. de Jager et al., 1988; H. Nieuwenhuijzen & C. de Jager, 1990) that are still used in state-of-the-art stellar evolution codes to model the RSG phase of massive stars (N. Mauron & E. Josselin, 2011; P. Massey et al., 2021).
@@ 2 Rationale of this study
The motivation behind this work is to address (a) whether a single star may incur heavy mass loss in the final stages before it explodes, that can be correlated with the presence of a dense CSM near the star, and (b) how the progenitor mass impacts the mass loss rates, CSM densities, and wind velocities. In this regard, our study accounts for progenitors of Type II supernovae, and we compare the CSM profiles of well-observed interacting SNe such as SN 2023ixf (J. E. Jencson et al., 2023; W. V. Jacobson-Galán et al., 2023; E. A. Zimmerman et al., 2024; N. Smith et al., 2023; K. A. Bostroem et al., 2024; N. A. J. et al., 2025), SN 2005ip (N. Smith et al., 2009; S. Katsuda et al., 2014), SN 1998S (N. Chugai et al., 2002; M. Pozzo et al., 2004; C. Fransson et al., 2005; J. Mauerhan & N. Smith, 2012; I. Shivvers et al., 2015), SN 2017hcc (P. Chandra et al., 2022), SN 2020ywx (R. Baer-Way et al., 2025). We argue that for single stars, enhanced pulsation near the end of their lives provides conditions favorable for escalated mass loss (S.-C. Yoon & M. Cantiello, 2010; Moriya, Takashi J. & Langer, Norbert, 2015). To demonstrate this effect, we simulate the evolution of massive stars with initial masses between $12$ and 2020 M, from their main-sequence until the end of neon burning (or until the end of core O-burning, in some cases) using the MESA code, applying advanced mass loss prescriptions to incorporate effects of radial pulsations in the outer envelope of RSGs (Section 4).
@@ 3 Mass loss in RSGs
Mass loss from RSGs is a fundamental yet still uncertain ingredient in the late stages of stellar evolution (H. J. Lamers & E. M. Levesque, 2017; R. M. Humphreys, 2025; S. D. Van Dyk, 2025; J. T. van Loon, 2025; S. Ekström & C. Georgy, 2025). The physical mechanisms responsible for this process directly shape the CSM, influencing both the pre-SN environment and the post-explosion observables, including light curves, spectra, and dust formation (E. M. Levesque, 2017). Observations of Type II SNe with dense CSM (e.g., SN 2009ip, SN 2010jl, SN 2023ixf and so on) have revealed that enhanced and often eruptive mass loss can occur shortly before CC (T. Zhang et al., 2012; E. O. Ofek et al., 2013; K. A. Chugai et al., 2023). In this section, we describe potential driving mechanisms for steady line-driven winds, pulsation-driven outflows including dynamical ejections and/or eruptions, turbulent pressure and convective shocks creating extended atmosphere/chromosphere for dust-driven winds, and enhanced mass loss due to binary interaction for RSGs.
@@ 3.1 Line-driven wind
Line-driven winds are driven by the transfer of momentum from photons to metal lines in the stellar atmosphere. This mechanism is efficient in hot, compact stars e.g., OB stars (J. S. Vink, 2017), but is considerably less effective in RSGs due to their cooler photospheres ($T_{\mathrm{eff}}\lesssim 4000$ K), where fewer atoms are ionized and UV photons are scarce (J. I. Castor et al., 1975; L. B. van den Hoek & M. A. T. Groenewegen, 1997). Typical mass loss rates from line-driven winds in RSGs range from $\sim 10^{-7}$ to $10^{-6}$ M $\rm{yr}^{-1}$ (E. R. Beasor et al., 2021; J. Merritt et al., 2025), which is inadequate to explain the dense CSM observed in the majority (upto $80\%$) of all Type II SNe with early-time observations (V. Morozova et al., 2018; F. Förster et al., 2018).
@@ 3.2 Pulsation driven mass loss
Many RSGs exhibit large-amplitude radial pulsations with periods of several hundred days (M. D. Soraisam et al., 2018; Y. Ren et al., 2019), as observed in long-period variables (LPVs). Such radial motions in the H-rich envelope of RSGs are driven by the so-called $\kappa$-mechanism (N. Baker & R. Kippenhahn, 1965; T. Gastine & B. Dintrans, 2008) - an interplay between excess radiation pressure and gravitational pull on the outer layers of the RSG envelope. These pulsations can grow in amplitude to produce shocks at the stellar surface that levitate material above the photosphere, reducing the effective gravity and allowing other processes (e.g., radiation pressure on dust and molecules) to accelerate the material outward. However, the exact mechanism of mass loss due to the complex interplay of pulsations and convection in RSG envelopes remains ill-understood (T. Gastine & B. Dintrans, 2010), and attempts to reduce these inherently three-dimensional phenomena into one-dimensional prescriptions for use in evolutionary codes remain limited (S.-C. Yoon & M. Cantiello, 2010; Moriya, Takashi J. & Langer, Norbert, 2015; V. A. Bronner et al., 2025).
@@ 3.2.1 Dynamical Ejections
In advanced nuclear burning stages, vigorous convection and shell flashes can lead to hydrodynamical instabilities and formation of shock waves in the outer envelopes of RSGs. If these internal shocks propagate outward and reach the stellar surface, they can trigger impulsive ejections of envelope material. These “dynamical ejections” are thought to be responsible for the formation of dense, compact shells detected in SN progenitors such as SN 2009ip (J. C. Mauerhan et al., 2013; N. Smith, 2014). Simulations suggest that mass loss rates in such events can temporarily exceed $10^{-2}$ M $\rm{yr}^{-1}$, lasting sometimes for days to weeks or even sustain repeatedly for centuries (M. Clayton, 2018).
@@ 3.2.2 Eruptive mass loss
Eruptive mass loss may occur through surface instabilities driven by opacity changes, pulsationally enhanced convective transport, wave-driven energy deposition or super-Eddington luminosities in RSG envelopes (E. Quataert & J. Shiode, 2012; J. Fuller, 2017; S. J. Cheng et al., 2024). These eruptions resemble luminous blue variable (LBV) behavior, though they are seen in cooler RSGs as well. The pre-SN outbursts seen in light echoes (e.g., SN 2013cu) are likely signatures of such events, with inferred $\dot{M}\sim 10^{-4}$$10^{-2}$ M $\rm{yr}^{-1}$ over timescales of months to years (G. Gräfener & J. S. Vink, 2016). Furthermore, statistical analysis of luminous stars in the Milky Way (C. S. Kochanek, 2011) and the study of $\eta$ Carinae analogs in nearby galaxies (R. Khan et al., 2015) reveal the presence of massive shells which hint towards instabilities triggered at the onset of C-burning leading to eruptions (R. Margutti et al., 2017).
@@ 4.1 Pulsation-Driven Superwind (PDSW)
Red Supergiants (RSGs) often exhibit long-period variability (M. D. Soraisam et al., 2018; Y. Ren et al., 2019; B. Hsu et al., 2024) that has been associated with changes in stellar radii due to excitation of different modes of pulsational instabilities through the $\kappa$-mechanism (H. Saio et al., 2016). Theoretically, A. Heger et al. (1997) demonstrated the growth of such pulsational instabilities in hydrodynamical evolutionary calculations with sufficient temporal resolution and found that the growth rates in their non-linear numerical calculations to be proportional to the luminosity ($\mathrm{L}$) to mass ($\mathrm{M}$) ratio. S.-C. Yoon & M. Cantiello (2010) showed that the growth rate ($\eta$) of the amplitude of the surface velocities increases with $L/M$ ratio and/or decreasing thermal (Kelvin-Helmholtz) time scale of the envelope ($\tau_{KH,env}$), according to the relation:
@@
where $R$ is the stellar radius, $\tau_{KH,env}:=GMM_{env}/LR$ ($M_{env}$ is the mass of the H-rich envelope ) and $C_{1}=9.219\times 10^{-6},~C_{2}=0.0393$ which are calibrated against their hydrodynamic models evolved until shock breakout at the surface. Figure 1 shows a linear least-squares fit for the relation given by Equation 1 for hydrodynamic MESA model sequences with initial masses between 121812-18 M evolved with timesteps $\delta t\leq 0.01~\text{years}$ until shocks develop at the stellar surface, which gives $C_{1}=9.393\times 10^{-6},C_{2}=0.0797$, in close agreement with the values of S.-C. Yoon & M. Cantiello (2010).
Figure 1: Growth rate of radial oscillations, $\eta$, as a function of stellar parameters (radius, mass, and Kelvin-Helmholtz time-scale of the envelope); the solid line represents a linear least-squares fit.
Figure 2: (Top) Evolution of the stellar radius (black curve plotted on the left axis) in a set of hydrodynamic MESA model sequences with initial mass $M_{\mathrm{init}}=15~\mathrm{M}_{\odot}$ in the RSG phase until shock breakout at the surface when Mach Number ($v_{surf}/c_{sound}$) exceeds unity (shown in red on the right axis). (Bottom) Corresponding evolution of a set of MESA models with varying initial masses between $12-18\ M_{\odot}$ in the HR diagram.
The top panel of Figure 2 shows a typical hydrodynamical model for a MESA run with initial mass $M_{\mathrm{init}}=15$ M illustrating the growth of the radial pulsations until shock break-out at the surface where the Mach number reaches unity. The bottom panel of Figure 2 shows the corresponding evolution of a set of MESA runs with varying initial masses between 121812-18 M in the Hertzspring-Russell (HR) diagram. Capturing these radial pulsations with the typical time period, $P\sim 2-5~yrs$, requires using much smaller time steps (B. Paxton et al., 2013), that makes modeling such pulsations a challenging task for evolutionary codes which typically use much larger time steps during the onset and growth of these pulsations (B. Paxton et al., 2013; V. A. Bronner et al., 2025). We do not attempt to resolve these pulsations for our full evolutionary runs (from pre main-sequence until core neon/oxygen depletion) for computational feasibility but incorporate potential impact of the growth of such pulsations observed in advanced phases (beyond core-helium burning) on the mass loss history of our stellar models in a phenomenological approach, similar to that of S.-C. Yoon & M. Cantiello (2010). To do this, we evolve a set of hydrostatic MESA models until end of core-helium burning using the standard Dutch wind scheme in MESA, scaled down by a factor of $0.2$ for reasons explained in Section 4. Thereafter, each run is continued with the hydrodynamic term on and replacing the wind mass loss rates with the ”superwind” (SW) mass loss prescription of S.-C. Yoon & M. Cantiello (2010) given by
@@
for $\eta>1$ and $T_{\rm{eff}}<4000~\rm{K}$, using the growth rate expression as given in Equation 1 with our calibrated constants. Here, $\dot{M}_{\mathrm{JNH88}}$ is the mass loss rate from de C. de Jager et al. (1988) scaled down by a factor of $0.2$, and $\alpha$ is a free parameter which sets the strength and consequently the duration of the SW-phase (Section 4.2). Unlike S.-C. Yoon & M. Cantiello (2010) who used a fixed $\alpha~=~1.75$, we conduct a thorough exploration of the dependence of the SW on both the initial mass $\mathrm{M}_{init}=12-20$ M and value of $\alpha=0-5$, motivated by observational evidence for enhanced mass loss in late stage of evolution of RSG progenitors of Type II SNe.